For the motion of a particle, velocityย  'v' depends on displacement x as v=20/(3x-2).If at t=0,x=0 then at what time t, the x=20?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sindhuja

 v=20/(3x-2)

v= dx/dt=20/(3x-2)

 (3x-2)dx = 20 dt

(3x-2)dx = 20 dt

intigrate

 3 x2/2 -2x  =20t +C

given at t=0 ,x=0

so C=0

3 x2/2 -2x  =20t

now put x =20t

3*202/2 -2*20 =20t

or t= 30-2 =28 sec

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin